# Maintaining Adversarial Robustness in Continuous Learning

### v1

@@ Abstract
Adversarial robustness is essential for security and reliability of machine learning systems. However, the adversarial robustness gained by sophisticated defense algorithms is easily erased as the neural network evolves to learn new tasks. This vulnerability can be addressed by fostering a novel capability for neural networks, termed continual robust learning, which focuses on both the (classification) performance and adversarial robustness on previous tasks during continuous learning. To achieve continuous robust learning, we propose an approach called Double Gradient Projection that projects the gradients for weight updates orthogonally onto two crucial subspaces – one for stabilizing the smoothed sample gradients and another for stabilizing the final outputs of the neural network. The experimental results on four benchmarks demonstrate that the proposed approach effectively maintains continuous robustness against strong adversarial attacks, outperforming the baselines formed by combining the existing defense strategies and continual learning methods.
@@ 1 Introduction
Continual learning and adversarial robustness are distinct and important research directions in the machine learning (ML) community, each of which has witnessed significant advances. The former addresses a critical challenge known as catastrophic forgetting, where a neural network trained on a new task typically exhibits a dramatic drop in its performance on previously learned tasks if the model cannot revisit the previous data [1]. The latter focuses on developing defenses against adversarial attacks that can deceive models into confidently misclassifying objects by adding subtle targeted perturbations to the input images often imperceptible to human observers [2].
However, the evolution of the neural network’s adversarial robustness in context of continuous learning remains unexplored. Meanwhile, in our experiments, we observe that adversarial robustness gained by well-designed defense algorithms suffers from a phenomenon similar to catastrophic forgetting. Just as in previous work when a dramatic performance drop was observed during continuous learning, an intriguing challenge arises: how can we maintain the adversarial robustness on previous tasks? In other words, we expand the objective in continuous learning to concurrently encompass performance and adversarial robustness.
In this paper, we present a solution inspired by Input Gradients Regularization (IGR) [3] and Gradient Projection Memory (GPM) [4]. With the hypothesis that a neural network with smooth gradients of original samples are more robust against adversarial examples, IGR optimizes neural networks by adding to the objective function a regularization term to penalize gradients of the classification loss with respect to the original samples, which ensures that the predictions of neural networks will not change significantly if any original sample changes locally. At the same time, GPM constraints the back-propagation gradients (used for updating neural network parameters) into the orthogonal direction to the crucial gradient subspace of previous tasks, of which base vectors are formed by performing singular value decomposition on layer-wise inputs of the neural network. GPM can effectively stabilize the final outputs of a neural network, guaranteeing consistent prediction results for data from previous tasks, even after adjusting its weights to accommodate new tasks.
Specifically, we hypothesize that, for a neural network that has smoothed the gradients of samples and is therefore robust for the current task, this robustness will hold if we stabilize these smoothed gradients during continuous learning. Compared to GPM, the crucial gradient subspaces in our approach consist of two parts: one exists in the layer-wise inputs, and the other exists in gradients of these inputs with respect to samples. When learning new tasks, the back-propagation gradients are firstly projected perpendicular to the basis vectors that span the subspaces before being used for weight updates. Our contributions are summarized as follows:
@@
    We introduce a novel capability, referred to as continual robust learning, which is valuable for the adaptive evolution of neural networks across multiple tasks. The key distinction between this and traditional continual learning lies in its focus on not only performance but also adversarial robustness.
@@
    We propose the Double Gradient Projection (DGP) approach, which facilitates continual robust learning by orthogonal projection of back-propagation gradients onto two crucial subspaces, stabilizing both sample gradients and final outputs across previous tasks.
@@
    We demonstrate the effectiveness of our approach on four benchmarks, showcasing its superiority over the baselines established by combining existing algorithms solely for adversarial robustness or continual learning.
@@ 2 Background
In this section, we introduce the preliminary concepts underlying our work.
2.1 Sample Gradient Regularization
The robustness of the neural network trained with IGR has been demonstrated in adversarial examples across multiple attacks, architectures and datasets [3]. IGR optimizes a neural network $f_{\mathbf{w}}$ by minimizing both the classification loss and the rate of change of that loss with respect to samples, formulated as:
@@ 2.1 Sample Gradient Regularization
2.2 Gradient Projection Memory
@@
where $\Delta\mathbf{W}_{p,t}^{l}$ denotes the change of weights in task $\mathcal{T}_{t}$ relative to task $\mathcal{T}_{p}$. Assuming $\mathbf{X}_{p,t}^{l}=\mathbf{X}_{p,p}^{l}$, a sufficient condition to guarantee $\mathbf{X}_{p,t}^{l+1}=\mathbf{X}_{p,p}^{l+1}$ is by imposing a constraint on $\Delta\mathbf{W}_{p,t}^{l}$ as [5].
@@
GPM [4] performs singular value decomposition (SVD) on $\mathbf{X}_{p,p}^{l}\in\mathbb{R}^{n\times m_{l}}$, where $n$ is the number of samples randomly drawn from the task $\mathcal{T}_{p}$ and $m_{l}$ is the number of features in an input of $l$-th layer:
@@
where $\alpha^{l}$ is a given threshold representing the trade-off between learning plasticity and memory stability of the model [6]. By establishing a dedicated pool $\mathcal{P}^{l}$ to retain base vectors $\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$ from previous tasks, GPM enforces the orthogonality of gradients with respect to these base vectors in the learning process of a new task:
@@ 3 Method
In this section, we propose an algorithm drawing inspiration from the IGR and GPM to effectively address the challenge of maintaining adversarial robustness in the continuous learning scenario, where revisiting previous data is not feasible. We hypothesize that adversarial robustness, which has been enhanced by IGR in previous tasks, will hold even after learning a sequence of new tasks, if we can stabilize the smoothed gradients of samples from previous tasks.
@@ 3.1 Constraint on Weight Updates
By applying the chain rule for derivatives of composite functions, the gradient of the model’s (with $L$ blocks) final output $\hat{\mathbf{y}}$ with respect to a sample $\mathbf{x}$ can be expressed in terms of recursive multiplication:
@@
$\begin{bmatrix}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{1}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(2\right)}}{\partial x_{m_{1}}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{m_{1}}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{L}^{l}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{1}}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{1}}}\end{bmatrix},$ (9)
@@ 3.1.1 Linear block
We then impose a constraint on weight updates, akin to Eq. 3 but specifically for stabilizing sample gradients (the core idea of this paper):
@@
$\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{i}}&\cdots\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{i}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{L}^{l}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{1}}{\partial x_{i}}&\cdots&\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{c}}{\partial x_{i}}\end{bmatrix}.$ (18)
According to Eq. 18, $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ transforms to a vector $\in\mathbb{R}^{m_{l}}$. This modification significantly reduces the computational time required for performing SVD on matrix $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times m_{l}}$, while relaxes the stringent guarantee for stabilizing $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}$ to a less restrictive one (see right-hand side of Eq. 9 and Eq. 18). The target of the constraint in Eq. 17 is altered from stabilizing the gradient of each final output with respect to each feature in the sample, to stabilizing the sum of gradients of each final output with respect to all features in the sample. This weak guarantee is sufficient to yield desirable results in our experiments for both fully-connected and convolutional neural networks.
@@ 3.1.2 Convolutional block
There are two differences between the convolution layer and the linear layer. First, the $\widetilde{\mathbf{W}}^{l}\in\mathbb{R}^{\left(c_{l}h_{l}\omega_{l}\right)\times\left(c_{l+1}h_{l+1}\omega_{l+1}\right)}$ is distinct from the weight matrix $\mathbf{W}^{l}\in\mathbb{R}^{\left(c_{l}k_{l}k_{l}\right)\times c_{l+1}}$, where each column represents a flattened convolution kernel. Here, $c_{l+1}$ ($k_{l}$) denotes the number (size) of kernels in the $l$-th layer, and $h_{l}$ ($\omega_{l}$) denotes the height (width) of the input $\mathbf{x}^{l}$. We give a simple example to illustrate the composition of $\widetilde{\mathbf{W}}^{l}$ in Appendix Fig. 1. The $\widetilde{\mathbf{W}}^{l}$ is sparse, with non-zero elements only present at specific positions of each column, corresponding to the input features that interact with a convolution kernel. To facilitate implementation, we input $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ into the $l$-th convolution layer after reshaping it from $\left(c_{l}h_{l}\omega_{l},\right)$ into $\left(c_{l},h_{l},\omega_{l}\right)$, and subsequently reshape the output from $\left(c_{l+1},h_{l+1},\omega_{l+1}\right)$ back to $\left(c_{l+1}h_{l+1}\omega_{l+1},\right)$, thereby obviating the intricate construction of $\widetilde{\mathbf{W}}^{l}$.
Second, the base vectors formed by directly performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times\left(c_{l}h_{l}\omega_{l}\right)}$ can be used to constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$ rather than $\Delta\mathbf{W}^{l}$. Consequently, we perform SVD after reshaping $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ into a matrix $\in\mathbb{R}^{\left(nh_{l+1}\omega_{l+1}\right)\times\left(c_{l}k_{l}k_{l}\right)}$ (see Appendix Fig. 2 for details).

### v2

@@ Abstract
Adversarial robustness is essential for security and reliability of machine learning systems. However, adversarial robustness enhanced by defense algorithms is easily erased as the neural network’s weights update to learn new tasks. To address this vulnerability, it is essential to improve the capability of neural networks in terms of robust continual learning. Specially, we propose a novel gradient projection technique that effectively stabilizes sample gradients from previous data by orthogonally projecting back-propagation gradients onto a crucial subspace before using them for weight updates. This technique can maintaining robustness by collaborating with a class of defense algorithms through sample gradient smoothing. The experimental results on four benchmarks including Split-CIFAR100 and Split-miniImageNet, demonstrate that the superiority of the proposed approach in mitigating rapidly degradation of robustness during continual learning even when facing strong adversarial attacks.
@@ 1 Introduction
Continual learning and adversarial robustness are distinct and important research directions in artificial intelligence, each of which has witnessed significant advances. The former addresses a critical challenge known as catastrophic forgetting, where a neural network trained on a sequential of new tasks typically exhibits a dramatic drop in its performance on previously learned tasks if the model cannot revisit the previous data [1]. The latter focuses on developing defenses against adversarial attacks that can deceive models into confidently misclassifying objects by adding subtle targeted perturbations to the input images often imperceptible to human observers [2].
However, the evolution of the neural network’s adversarial robustness in context of continuous learning remains underexplored. In our experiments, we observe that adversarial robustness enhanced by well-designed defense algorithms on previous data is easily lost when the neural network updates its weights to accommodate new tasks, resulting in a phenomenon similar to catastrophic forgetting. This presents an intriguing challenge: how can we maintain the adversarial robustness during continuous learning? In other words, the objective of continuous learning expands to concurrently encompass (classification) performance and adversarial robustness.
In this paper, we present a solution by proposing a novel gradient projection technique called Double Gradient Projection (DGP), which inherently enables collaboration with a class of defense algorithms that enhance robustness through sample gradient smoothing. DGP is grounded on a theoretical hypothesis that a neural network’s robustness can be maintained if the smoothness of sample gradients from previous data remain unchanged after weight updates. Specifically, when learning a new task, DGP projects the back-propagation gradients into the orthogonal direction to a crucial subspace before utilizing them for weight updates. This gradient subspace consists of two sets of base vectors derived from previous tasks, which are obtained by performing singular value decomposition on the layer-wise outputs of the neural network and the gradients of layer-wise outputs with respect to samples, respectively. Our contributions are summarized as follows:
@@
    We introduce the problem of robust continual learning in the scenario where data from previous tasks cannot be revisited.
@@
    We propose the Double Gradient Projection approach that stabilizes the sample gradients from previous tasks by orthogonally constraining the direction of weight updates. It can maintain robustness by collaborating with a class of defense algorithms that enhance robustness through sample gradient smoothing.
@@
    We validate the superiority of our approach on four image benchmarks. Furthermore, the experiment results indicate that without a tailored design, direct combination of existing continual learning and defense algorithms into the training procedure can be conflicting, resulting that the efficacy of the former is seriously weakened.
@@ 2 Background
In this section, we introduce the preliminary concepts underlying our work, including sample gradient smoothing and gradient projection.
2.1 Sample Gradient Smoothing
Input gradient regularization (IGR) [3]. The robustness of the neural network trained with IGR has been demonstrated across multiple attacks, architectures and datasets. IGR optimizes a neural network $f_{\mathbf{w}}$ by minimizing both the classification loss and the rate of change of that loss with respect to samples, formulated as:
@@ 2.1 Sample Gradient Smoothing
Adversarial training (AT) [4]. AT enhances robustness by incorporating adversarial examples generated by Fast Gradient Sign Method (FGSM)  [5] into training data. Compared to IGR, which explicitly smooths sample gradients by adding a regularization term into the loss function, AT achieves gradient smoothing implicitly.
2.2 Gradient Projection
@@
where $\Delta\mathbf{W}_{p,t}^{l}$ denotes the change of weights in task $\mathcal{T}_{t}$ relative to task $\mathcal{T}_{p}$. Assuming $\mathbf{X}_{p,t}^{l}=\mathbf{X}_{p,p}^{l}$, a sufficient condition to guarantee $\mathbf{X}_{p,t}^{l+1}=\mathbf{X}_{p,p}^{l+1}$ is by imposing a constraint on $\Delta\mathbf{W}_{p,t}^{l}$ as [6, 7]
@@
Gradient Projection Memory (GPM) [6], designed for improving continual learning ability, performs singular value decomposition (SVD) on $\mathbf{X}_{p,p}^{l}\in\mathbb{R}^{n\times m_{l}}$, where $n$ is the number of samples randomly drawn from the task $\mathcal{T}_{p}$ and $m_{l}$ is the number of features in an input of $l$-th layer:
@@
where $\alpha^{l}$ is a given threshold representing the trade-off between learning plasticity and memory stability of the neural network [8]. By establishing a dedicated pool $\mathcal{P}^{l}$ to retain base vectors $\left(\mathbf{V}^{l}_{k}\right)^{\mathrm{T}}$ from previous tasks, GPM enforces the orthogonality of gradients with respect to these base vectors in the learning process of a new task:
@@ 3 Method
In this section, we propose a novel gradient projection technique inspired from GPM to tackle the challenge of maintaining adversarial robustness in a continuous learning scenario, where revisiting previous data is not feasible. We hypothesize that if we can stabilize the sample gradients smoothed by defense algorithms such as IGR and AT on previous tasks, the adversarial robustness of the neural network will hold even after its weights update for learning a sequence of new tasks.
@@ 3.1 Constraint on Weight Updates
By applying the chain rule for derivatives of composite functions, the gradient of the neural network’s (with $L$ blocks) final output $\hat{\mathbf{y}}$ with respect to a sample $\mathbf{x}$ can be expressed in terms of recursive multiplication:
@@
$\begin{bmatrix}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{1}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(2\right)}}{\partial x_{m_{1}}}&\cdots&\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{m_{1}}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}^{L}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{1}}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{1}}}\end{bmatrix},$ (9)
@@ 3.1.1 Linear block
We then impose a constraint on weight updates for stabilizing sample gradients (the core idea of this work):
@@
$\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{1}^{\left(2\right)}}{\partial x_{i}}&\cdots\sum\limits_{i=1}^{m_{1}}\frac{\partial x_{m_{2}}^{\left(2\right)}}{\partial x_{i}}\end{bmatrix}\begin{bmatrix}\frac{\partial x_{1}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{1}^{\left(2\right)}}\\ \vdots&\vdots&\vdots\\ \frac{\partial x_{1}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}&\cdots&\frac{\partial x_{m_{3}}^{\left(3\right)}}{\partial x_{m_{2}}^{\left(2\right)}}\end{bmatrix}\cdots\\ \begin{bmatrix}\frac{\partial\hat{y}_{1}}{\partial x_{1}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{1}^{L}}\\ \vdots&\vdots&\vdots\\ \frac{\partial\hat{y}_{1}}{\partial x_{m_{L}}^{L}}&\cdots&\frac{\partial\hat{y}_{c}}{\partial x_{m_{L}}^{L}}\end{bmatrix}=\begin{bmatrix}\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{1}}{\partial x_{i}}&\cdots&\sum\limits_{i=1}^{m_{1}}\frac{\partial\hat{y}_{c}}{\partial x_{i}}\end{bmatrix}.$ (12)
According to Eq. 12, $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ transforms to a vector within the space $\mathbb{R}^{m_{l}}$. This modification significantly reduces the computational time required for performing SVD on matrix $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times m_{l}}$, while relaxes the stringent guarantee for stabilizing $\frac{\partial\hat{\mathbf{y}}}{\partial\mathbf{x}}$ to a less restrictive one (see right-hand side of Eq. 9 and Eq. 12). The target of the constraint in Eq. 11 is altered from stabilizing the gradient of each final output with respect to each feature in the sample, to stabilizing the sum of gradients of each final output with respect to all features in the sample. This weak guarantee is sufficient to yield desirable results in our experiments for both fully-connected and convolutional neural networks.
@@ 3.1.2 Convolutional block
There are two differences between the convolution layer and the linear layer. First, the $\widetilde{\mathbf{W}}^{l}\in\mathbb{R}^{\left(c_{l}h_{l}\omega_{l}\right)\times\left(c_{l+1}h_{l+1}\omega_{l+1}\right)}$ is distinct from the weight matrix $\mathbf{W}^{l}\in\mathbb{R}^{\left(c_{l}k_{l}k_{l}\right)\times c_{l+1}}$, where each column represents a flattened convolution kernel. Here, $c_{l+1}$ ($k_{l}$) denotes the number (size) of kernels in the $l$-th layer, and $h_{l}$ ($\omega_{l}$) denotes the height (width) of the input $\mathbf{x}^{l}$. We give a simple example to illustrate the composition of $\widetilde{\mathbf{W}}^{l}$ in Appendix Fig. 6. The $\widetilde{\mathbf{W}}^{l}$ is sparse, with non-zero elements only present at specific positions of each column, corresponding to the input features that interact with a convolution kernel. To circumvent the intricate construction of matrix $\widetilde{\mathbf{W}}^{l}$, we identify an alternative approach for implementing $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}^{l}}$: reshaping $\frac{\partial\mathbf{x}^{l}}{\partial\mathbf{x}}$ from $\left(c_{l}h_{l}\omega_{l},\right)$ to $\left(c_{l},h_{l},\omega_{l}\right)$, feeding it into the $l$-th convolution layer, and subsequently reshaping the output from $\left(c_{l+1},h_{l+1},\omega_{l+1}\right)$ back to $\left(c_{l+1}h_{l+1}\omega_{l+1},\right)$, i.e., $\frac{\partial\mathbf{x}^{l+1}}{\partial\mathbf{x}}$.
Second, the base vectors formed by performing SVD on $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}\in\mathbb{R}^{n\times\left(c_{l}h_{l}\omega_{l}\right)}$ (computed through Eq. 12), can be used directly to constrain the updates $\Delta\widetilde{\mathbf{W}}^{l}$ rather than $\Delta\mathbf{W}^{l}$ (see Appendix Fig. 7 for details). Consequently, we perform SVD after reshaping $\frac{\partial\mathbf{X}^{l}}{\partial\mathbf{X}}$ into a matrix $\in\mathbb{R}^{\left(nh_{l+1}\omega_{l+1}\right)\times\left(c_{l}k_{l}k_{l}\right)}$. This results that the base vectors have the same shape $\left(c_{l}k_{l}k_{l},\right)$ with the flattened convolution kernels in the $l$-th layer, and can be used directly to constrain their weight updates.